The perimeter of a rectangular field is 298 yards. If the width of the field is 64 yards, what is its length?

Answer by checkley79(3341)

P=2L+2W
298=2L+2*64
298=2L+128
2L=298-128
2L=170
L=170/2
X=85 YDS. FOR THE LENGTH.
PROOF:
298=2*85+2*64
298=170+128
298=298